Oil Wells offers 5.6 percent coupon bonds with annual payments and a yield to maturity of 6.94 percent. The bonds mature in seven years. The face value is $1,000. What is the current yield?
Answer #1

Annual coupon=1000*5.6%=56

Hence current price=Annual coupon*Present value of annuity factor(6.94%,7)+$1000*Present value of discounting factor(6.94%,7)

=56*5.40058109+1000*0.625199672

=$927.63(Approx)

Hence current yield=Annual coupon/Current price

=56/927.63

=6.04%(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=56[1-(1.0694)^-7]/0.0694

=56*5.40058109

2.Present value of discounting factor=1000/1.0694^7

=1000*0.625199672